OpenStudy (arindameducationusc):

New Trigonometry tutorial 1.2

OpenStudy (arindameducationusc):

\(\Huge\color{red}{Trigonometry} \) \(\Large\color{blue}{Trigonometry~ratios~of~the~sum~and~difference} \) \(\Large\color{blue}{of~two~angles} \) \(\Large\color{blue}{1)} \) \[\sin (A+B)=sinA cosB+cosAsinB\]

OpenStudy (arindameducationusc):

\(\Large\color{blue}{2)} \) \[\cos (A+B)=cosAcosB-sinAsinB\]

OpenStudy (arindameducationusc):

\(\Large\color{blue}{3)} \) \[\sin(A-B)=sinAcosB-cosAsinB\]

OpenStudy (arindameducationusc):

\(\Large\color{blue}{4)} \) \[\cos(A-B)=cosAcosB+sinAsinB\]

OpenStudy (arindameducationusc):

ya, please see the previous tutorials by me and @rvc before seeing this one..

OpenStudy (zzr0ck3r):

I think we should have a tutorial section.

OpenStudy (arindameducationusc):

\(\Large\color{blue}{5)} \) \[\sin C+sinD=2\sin \frac{ C+D }{ 2}\cos \frac{ C-D }{ 2 }\]

OpenStudy (arindameducationusc):

\(\Large\color{blue}{6)} \) \[sinC-sinD=2\cos \frac{ C+D }{ 2 }\sin \frac{ C-D }{ 2 }\]

OpenStudy (arindameducationusc):

\(\Large\color{blue}{7)} \) \[cosC+cosD=2\cos \frac{ C+D }{ 2}\cos \frac{ C-D }{ 2 }\]

OpenStudy (arindameducationusc):

\(\Large\color{blue}{7)} \) \[cosC-cosD=2\sin \frac{ C+D }{ 2 }\sin \frac{ D-C }{ 2 }\]

OpenStudy (arindameducationusc):

sorry previous was \(\Large\color{blue}{8} \)

OpenStudy (sepeario):

very extensive and useful tutorial, ty!

OpenStudy (arindameducationusc):

Thank you @Sepeario I will prove some of the above tomorrow, don't worry. I will have to take rest, tomorrow I am running a marathon

OpenStudy (sepeario):

haha nice!

rvc (rvc):

good job !

OpenStudy (arindameducationusc):

\(\Large\color{red}{Proof~of=>} \) sin(A+B)=sinAcosB+cosAsinB

OpenStudy (arindameducationusc):



OpenStudy (arindameducationusc):

A didn't come in the figure... its right of Q